OpenStudy (anonymous):

Which of the following is closest to [integrand] (upper limit is 3 & lower limit is 2) ((x^4)/2 )- (4e^x) dx? A: -29.7 B: -18.3 C: 78.3

OpenStudy (anonymous):

I'm guessing we can use a calculator to evaluate the limit then.

OpenStudy (anonymous):

@Brutal It's an integral :-)

OpenStudy (anonymous):

Oh yea, I meant integral, those limits threw me off. @lanad8 the official notation would be: Int(f(x),x,a,b) where a is the lower limit, and b is the upper limit, so in this case: Int( ((x^4)/2)- (4e^x), x, 2, 3)

OpenStudy (anonymous):

I took the integral and found 1x^5/10 -4e^x. What did you get at least for the integral without evaluating it?

OpenStudy (anonymous):

Actually this is real easy to integrate: \[\int((x^4)/2- (4e^x))dx = \] \[(1/2)*\int x^4dx - 4\int e^xdx=\] \[(1/2)*(x^5)/5-4e^x+C=\] \[x^5/10-4e^x+C\]

OpenStudy (anonymous):

that's without evaluating the integral at the limits.

OpenStudy (anonymous):

I evaluated it from 2 to 3 and got -29.7

OpenStudy (anonymous):

ok, thanks brinethery & Brutal!

OpenStudy (anonymous):

yvm! May I recommend theintegralcalc's channel? She's good at throwing tough examples out :-)

OpenStudy (anonymous):

i'll check out her channel. thank you :)

OpenStudy (anonymous):

PatrickJMT rules on youtube.

OpenStudy (anonymous):

Hmm... she had the integral of 1/x^5. Wouldn't the antiderivative be ln|x^5|?

OpenStudy (anonymous):

First of all no, she had the integral of \[x^5\] Second of all, no the antiderivative of \[1/x^5\] is not \[ln|x^5|\] \[\int(1/x^5)dx=\int(x^{-5})dx=x^{-5+1}/(-5+1)+C \]

OpenStudy (anonymous):

Or: \[-x^{-4}/4+C\]

OpenStudy (anonymous):

But then what's the derivative of ln(x)?

OpenStudy (anonymous):

\[(d/dx)(ln(x))=(1/x)\]

OpenStudy (anonymous):

See?!

OpenStudy (anonymous):

lol :-)

OpenStudy (anonymous):

\[(d/dx)(ln(x^4))=(1/(x^4))*4x^3=4x^3/x^4=4/x\]

OpenStudy (anonymous):

The only time you use Ln is when you have \[1/x\]because if you try to integrate that the proper way, you'll get a 0 in the denominator.

OpenStudy (anonymous):

\[\int(1/x)dx=ln|x|+C\] you can't do this: \[\int(1/x)dx=\int(x^{-1})dx=x^{-1+1}/(-1+1)+C\]

OpenStudy (anonymous):

Oh I see what you're saying.

OpenStudy (anonymous):

See the bottom only gives you trouble with the power of -1. Otherwise you're good.

OpenStudy (anonymous):

It's been like... 6 years since I've taken calc so I've forgotten a lot hehe

OpenStudy (anonymous):

Thanks for explaining that :-)

OpenStudy (anonymous):

np